# Possible Unconventional Surface Superconductivity in the Half-Heusler YPtBi

### v1

@@ Abstract
We report an extensive extensive study of the noncentrosymmetric half-Heusler topological superconductor YPtBi, revealing unusual relation between bulk superconductivity and the appearance of surface superconductivity at temperatures up to 3 times the bulk transition temperature. Transport measurements confirmed the low carrier density of the material and its bulk superconducting transition, which was also observed in ac susceptibility through mutual inductance (MI) measurements. However, a weak signature of superconductivity in the MI measurements appeared much above the bulk transition temperature, which was further observed in scanning tunneling spectroscopy. Polar Kerr effect measurements suggest that while the bulk superconductor may exhibit an unusual nodal superconducting state, only the surface state breaks time reversal symmetry. Complementary tunneling measurements on LuPtBi are used to establish the observations on YPtBi, while density-functional theory (DFT) calculations may shed light on the origin of this unusual surface state.
Introduction:
@@
Topological superconductivity can be realized by inducing superconductivity at the surface of a three dimensional topological insulator (3D TI) [1]. This can be realized through interface engineering, or by using the surface states of a bulk superconductor [4, 3, 2]. These mechanisms generally lead to time-reversal invariant topological states [1], and their critical temperature is determined by that of the bulk superconductor. An alternative approach, which expands the limits of these techniques, is utilizing flat surface bands which can support superconductivity without a proximity effect. Such superconducting surface states have been observed in KZnBi [5] and t-PtBi [6]. However, the topological properties of these states, particularly as related to the origin and nature of their superconducting properties, are not fully understood.
@@
Here, we explore the possible formation of a superconducting surface state at the (111) surface of the half Heusler compound YPtBi through extensive measurements that include transport, ac susceptibility measurements through mutual inductance, scanning tunneling microscopy and spectroscopy and polar Kerr effect. While robust bulk superconductivity appears below , we find that superconductivity at the surface persists to temperatures . Furthermore, our Kerr effect measurements suggest that unlike the bulk, this superconducting surface state breaks time reversal symmetry (TRS), thus revealing a delicate interplay between the surface and bulk states in this paradigmatic topological system. We complement the scanning tunneling spectroscopy measurements on YPtBi with measurements LuPtBi to substantiate our arguments. To further support our experimental findings, we performed ab-initio calculations on the Bi, Lu and Pt terminations of the (111) direction, finding that a Bi-termination surface exhibits a van-Hove singularity (VHS) near the Fermi energy, which could be a source for enhanced susceptibility for a superconducting surface state. A recent theoretical study [7] explored a similar surface band structure, finding a possible domain for the onset of unconventional chiral topological surface superconductivity ahead of the bulk transition, in agreement with our results. Our results reveal a new platform for 2D chiral superconductors in which the topologically-protected surface state condense into a TRS-breaking superconducting state prior to the bulk.
@@
In a half-Heusler compound, three metallic elements are arranged in interpenetrating fcc lattices ( structure) [8]. The choice of elements, with thousands of possible combinations, enables control over a wide variety of properties in the resulting compounds: structures can be metallic, semiconducting, magnetic or non-magnetic, superconducting, and they can have topological surface states [9, 10]. Of these materials, YPtBi and LuPtBi have the largest inverted gaps [11], leading to topologically protected surface states [12]. Furthermore, YPtBi and LuPtBi are bulk superconductors, with critical temperatures of 0.8 K and 1 K respectively [13, 14]. The bulk superconductivity should be a combination of odd and even parity states, due to the lack of inversion symmetry [15]. While much work was done to explore the bulk superconductivity in these materials[13, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], studies of surface states and their propensity to unconventional superconductivity have been limited [7].
Methods:
Single crystals of LuPtBi and YPtBi were grown by self-flux solution growth method, with Bi used as the flux medium [29, 30, 31]. The composition and structure were verified using dispersive X-ray analysis and Laue X-ray diffraction, respectively. We used X-ray Photoelectron Spectroscopy (XPS) to characterize the elemental surface composition of the samples. With a typical measurement depth of about a nanometer, pristine crystals showed traces of carbon, oxygen, and bismuth on the surface. The carbon traces were eliminated after several seconds of Ar ion milling (equivalent to a few nm), and all three elements: Lu, Pt and Bi seem to appear with the correct 1:1:1 ratio, together with traces of oxygen. The surface composition did not change upon subsequent milling of several more nanometers of the surface, confirming that no other elements exist in the samples.
Resistivity and Hall effect on both types of samples, confirming the low carrier densities, /cm, and relatively high mobilities, typically cm/V-s at helium temperature. Sharp bulk superconductivity was verified with both, resistive transition and two-coils mutual inductance (MI) measurements.
Scanning Tunneling Microscopy (STM) and spectroscopy (STS) measurements were performed in a modified UNISOKU Co., USM1300SHeTM system with Ultra High Vacuum (UHV), 350 mK base temperature, and magnetic field capabilities. Our studies focused on samples that exhibited a clear (111) surface such as shown in the insets of Fig. 1. The data shown in this work were obtained on native as-grown surfaces without any ex-situ preparation. We performed in-situ surface cleaning by removing the top few surface layers with the STM tip itself. The details of this procedure are described in the Supplemental Material [31], and are based on a similar procedure previously used to expose superconducting regions in vacuum-cleaved YPtBi crystals [32].
High-resolution measurements of Polar Kerr Effect (PKE) were performed using a continuous wave, Zero-Area-Loop Sagnac Interferometer (ZALSI) [33, 34], in a variable temperature He system. This instrument was successfully used to detect TRSB-breaking order parameters in unconventional superconductors and subtle effects of magnetism in a variety of quantum materials (see Supplemental Information [31]).
Results and Discussion:
@@ Methods:
Figure 1 shows representative base temperature (350 mK) spectra of pristine surface and tip-cleaned surface for YPtBi and LuPtBi samples. While the tip-cleaned surface improved the spectra, particularly in reducing the observed zero bias conductance to very low values, the pristine surface spectra pointed to a similar large gap. At the same time the observed spectra are rounded, even well below temperatures corresponding to the observed large gap. Such a behavior could be because of nodal gap structure, high effective temperature due to RF heating or due to an intermediate thin metallic layer which depress the gap by proximity effect [35]. The fact that on occasion we observe a flat gap as in Fig. 1b suggests that the rounding of the gap is not due to nodes in the gap.
The large gaps observed at base temperature indicate that there is a superconducting system within either LuPtBi or YPtBi which involves a transition temperature higher than the bulk. Indeed, temperature dependence of tunneling spectra of both materials indicate gaps that persist much above the bulk . Focusing on YPtBi, we show in Fig. 2a a typical set of spectra taken at different temperatures. Similar data on YPtBiwas reported in [32]. Since the location of the coherence peaks is not a good measure of the size of the gap due to the rounded data at all temperatures, some assumptions are required to estimate the gap as a function of temperature. In a first approach we assume that “noise” interferes with the measurement and thus each spectrum is considered as a standard BCS spectrum taken at an effective temperature higher than the actual sample temperature . Such an approach is similar to assuming a an unknown broadening of the spectrum [36]. The fit to each spectrum yields the gap and the effective temperature . Alternatively we can estimate the gap from maximum slope of the spectrum before it turns into the coherence peak (this is similar to using the maximum derivative of the spectrum [37]). Both estimates give a maximum gap of 0.4 to 0.41 meV and the fit to the BCS gap function with yields a K.
@@
To further investigate the onset of superconductivity, we measured the ac susceptibility of both materials using a mutual inductance technique [38]. Both materials showed a large increase of the diamagnetic response at their respective bulk transition temperatures (Figure 3), in agreement with previous reports [13, 14]. However, a careful measurement of YPtBi revealed a small increase, 0.5% of the change due to the bulk transition, which onsets at 3 K (Fig. 3a). This may indicate that a small volume fraction of the sample turns superconducting at a higher temperature.
@@
The observation of a similar transition temperature for the vanishing of the superconducting gap in tunneling and a weak onset of inductive signal in mutual inductance points to an intrinsic origin of superconductivity that onsets at a temperature much above the bulk . We envision two possible origins for the above observations. First, given our inability to make clean, atomically flat cleaves that are known to expose the (111) plane, the signals may alternatively be attributed to another compound, created during the surface cleaning, a combination of the HH elements, possibly combined with contaminants, particularly oxygen. Superconducting compounds that may be formed this way are listed in Table 1. While the of both amorphous and polycrystalline Bi are suitable, such compounds on the surface would generate similar transitions for both LuPtBi and YPtBi, inconsistent with the huge difference between the two compounds observed in STM measurements. We further note that LuPtBi and YPtBi are very similar chemically and structurally (lattice constants Å and Å, respectively [11]). Thus, any strained inclusion of a different compound on the surface would be expected to yield a similar gap. Moreover, the similarity in the behavior of LuBi and YBi may rule this possibility out as well, particularly since they require pressure to exhibit superconductivity.
Compound | Bi | a-Bi | p-Bi | BiPt | BiPt | YBi | LuBi |
@@
The alternative explanation for the observed higher in the two compounds is the creation of a chiral surface state. Particularly, since both LuPtBi and YPtBi are known to be topological with a large band inversion [11], it was suggested that superconductivity associated with this material is topological and may break TRS [22, 23]. To test this possibility, particularly for YPtBi, we performed high resolution polar Kerr effect measurements using a zero-area-loop Sagnac interferometer (ZALSI) [33, 34]. By construction, the ZALSI is inherently reciprocal and thus by its symmetry it can distinguish between true TRSB Kerr effect and optical activity when reflected from a chiral medium. In our implementation we use two orthogonally counter propagating linearly polarized light beams, which are then converted into right and left circular polarizations using a quarter-wave plate. An objective lens focuses the light onto a small interaction region on the sample, and the reflected light beams are converted back to linear polarizations with exchanged polarization directions. Owing to the reciprocity of the apparatus, any non reciprocal phase shift will appear at the detector only if TRS is broken through the interaction of the two circularly polarized beams with the sample [33, 34] (see supplementary material [31]).
@@
We cooled down the sample to 300 mK from a high temperature in the presence of a small external magnetic field, and then removed the field and measured the Kerr signal while warming up the sample (Figure 4a). For small training fields ( Oe) the initial signal (at 300 mK) following the removal of the magnetic field was very small, consistent with zero (within our resolution of rad). However, increasing the temperature from base temperature, a finite Kerr signal of order rad appeared above the bulk . The signal was independent of the field strength and changed sign upon changing the direction of the training field (Figure 4b). These features indicate a TRS breaking state, and are generally similar to those observed in chiral superconductors [47, 48].
@@
With strong structural and electronic structure similarities to GdPtBi, we estimate the optical penetration depth of YPtBi at nm [49]. Thus, the Kerr signal can have both bulk and surface contributions. In particular above the bulk the observed Kerr signal may be a result of TRS breaking associated with the surface superconductivity - a favorable possibility since Kerr effect persists very close to the surface , or it may be associated with vestigial order [50] originating from the bulk superconducting state. However, the fact that Kerr effect is observed up to three times the bulk makes the vestigial order possibility unlikely.
@@
While low field training (i.e. training in Oe) yields similar onset and size of Kerr signal, which follows the direction of the training field, cooling the sample at higher magnetic field to base temperature (0.3 K), turning the field to zero and measuring while warming up the sample in zero magnetic field yields a signal that is zero within our resolution of rad. However, cooling the sample at higher magnetic field to base temperature (0.3 K), but on the way turning the magnetic field to zero at K, then measuring the Kerr effect at zero field from base temperature up yields results similar to the low field training experiments. Fig. 5 summarizes these findings.
@@
Such a peculiar effect may point to a delicate interplay between the vortex state that appears in the bulk material upon cooling in a field much below the bulk and the effect on the surface state. For example, cooling at lower field and removing the field at low temperatures may result in low vortex occupancy due to weak surface and geometrical barriers (note that Oe, while the thermodynamic critical field is Oe for the bulk superconducting state of these materials). Above a field of order Oe, enough vortices may remain in the bulk, which upon crossing the bulk Tc will move out of the sample, thus possibly “scrambling” the single domain nature of the TRS-breaking state of the surface superconductivity. We note that the apparent suppression of the Kerr effect at temperatures below suggests points to a TRS-preserving bulk superconducting state. This is consistent with the observations of nodes in the bulk and the analysis of the possible spin and orbital pairing states [23]. Thus, the magnetization of the surface state interacts with a conventional superconducting order parameter. This is similar to magnetic superconductors, where the formation and motion of vortices can change the ferromagnetic order parameter [51].
@@
Taken together, the STM, mutual inductance and PKE results indicate that the (111) surfaces of YPtBi and LuPtBi host a superconudcting state which breaks TRS. To understand the origin of this state, we studied the surface band structure of LuPtBi (see Supplementary Information [31]). In Bi-terminated (111) surfaces, we found a van-Hove signularity close to the -point, giving rise to an enhanced density of states near the Fermi level. This could be the driving force behind the enhanced surface . The calculated band structure of the Lu terminated surface, did not show van-Hove singularities. Schwemmer et al. [7] have recently reported a detailed study of the surface band structure and possible superconductivity of the Bi-terminated (111) surface of LuPtBi. While the gross features of their surface band structure are similar to our calculation, they also revealed flat-band regions near the Fermi energy, and calculated propensity to a superconducting instability. In particular, they argue that the transition temperature of the surface state is independent of the bulk transition, and exhibit vestigial time-reversal symmetry breaking in it fluctuation regime. Since inversion symmetry is broken, simple singlet and/or triplet states were dismissed and pairing was tested for pairs of electrons that are connected by the TR operator. The result has been a pair-amplitude that transforms under the E irreducible representation of the surface’s point group (relevant to the (111) crystal orientation) in the presence of strong Kane-Mele type and Rashba type spin-orbit coupling [52]. Furthermore, they found that a chiral superconducting state, which spontaneously breaks TRS, is energentically favorable.
@@
In conclusion, we presented above a comprehensive study of scanning tunneling spectroscopy of the tip-cleaned (111) surface of the half-Heusler alloys YPtBi and LuPtBi. While a bulk of 0.8 K is observed for both systems, we find a much larger corresponding gap in our tunneling studies with ’s of 2.3 K and 6.5 K for YPtBi and LuPtBi respectively. Mutual inductance measurements of YPtBi further revealed a small diamagnetic response that onsets around . Polar Kerr effect on YPtBi indicate that TRS is broken near or at . Together, our data indicate that chiral surface superconductivity may precede bulk condensation in these materials, a hypothesis which is strongly supported by recent calculations [7]. While additional work is required to determine the order parameter of these states, our results open a new direction to engineering topological superconducting states in half Heusler compounds. While our study was conducted on single crystals, Kim et al. [53] recently demonstrated epitaxially growth of YPtBi films, which can also be encapsulated by an alumina film, preventing degradation of the surface. This system is therefore highly promising for quantum device applications.

### v2

@@ Abstract
We report an extensive study of the noncentrosymmetric half-Heusler superconductor YPtBi, revealing an unusual relation between bulk superconductivity and the possible appearance of surface superconductivity on the (111) oriented surface, at temperatures up to 3 times the bulk transition temperature. Transport measurements confirmed the low carrier density of the material and its bulk superconducting transition, which was also observed in ac susceptibility through mutual inductance (MI) measurements. However, a weak signature of superconductivity in the MI measurements appeared much above the bulk transition temperature, which was further observed in scanning tunneling spectroscopy, pointing to a possible surface superconducting state. Polar Kerr effect measurements suggest that while the bulk superconductor may exhibit an unusual nodal superconducting state, only the surface state breaks time reversal symmetry. Complementary tunneling measurements on LuPtBi are used to establish the observations on YPtBi, while density-functional theory calculations may shed light on the origin of this unusual surface state.
@@
Topological superconductivity can be realized by inducing superconductivity at the surface of a three dimensional topological insulator [1]. This can be achieved through interface engineering, or by using the surface states of a bulk superconductor [4, 3, 2]. These mechanisms generally lead to time-reversal invariant topological states [1], and their critical temperature is determined by that of the bulk superconductor. An alternative approach, which expands the limits of these techniques, is utilizing flat surface bands which can support superconductivity without a proximity effect. Such superconducting surface states have been observed in KZnBi [5] and t-PtBi2 [6]. However, the properties of these states, particularly the origin and nature of their superconductivity, are not fully understood.
@@
Here, we explore the possible formation of a superconducting surface state at the (111) surface of the half Heusler compound YPtBi through extensive measurements that include transport, ac susceptibility measurements through mutual inductance, scanning tunneling microscopy and spectroscopy and polar Kerr effect. While robust bulk superconductivity appears below , we find that superconductivity at the (111) surface persists to temperatures . Furthermore, our Kerr effect measurements suggest that unlike the bulk, this superconducting surface state breaks time reversal symmetry (TRS), thus revealing a delicate interplay between the surface and bulk states. We complement these measurements with scanning tunneling spectroscopy and mutual inductance of (111) oriented LuPtBi to substantiate our arguments. To further support our experimental findings, we performed ab-initio calculations on the Bi, Lu and Pt terminations of the (111) direction, finding that a Bi-termination surface exhibits a van-Hove singularity (VHS) near the Fermi energy, which could be a source for for the superconducting surface state. A recent theoretical study [7] explored a similar surface band structure, finding a possible domain for the onset of chiral topological surface superconductivity ahead of the bulk transition, in agreement with our results. Our results suggest a new platform for 2D chiral superconductors in which a surface state condenses into a TRS-breaking superconducting state independently from the bulk.
@@
In a half-Heusler compound, three metallic elements are arranged in interpenetrating fcc lattices ( structure) [8]. The choice of elements, with thousands of possible combinations, enables control over various properties in the resulting compounds: structures can be metallic, semiconducting, magnetic or non-magnetic, superconducting, and they can have topological surface states [9, 10]. Of these materials, YPtBi and LuPtBi have the largest inverted gaps [11], leading to topologically protected surface states [12]. Furthermore, YPtBi and LuPtBi are bulk superconductors, with critical temperatures of 0.8 K and 1 K respectively [13, 14]. The bulk superconductivity should be a combination of odd and even parity states, due to the lack of inversion symmetry [15]. While much work was done to explore the bulk superconductivity in these materials [13, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], studies of the surface states and their propensity to unconventional superconductivity have been limited [7].
We studied single crystals of LuPtBi and YPtBi that were grown by self-flux solution growth method, with Bi used as the flux medium [31, 32, 33]. The crystal structure and composition were verified using dispersive X-ray analysis, Laue X-ray diffraction and X-ray Photoelectron Spectroscopy [33]. Transport measurements confirmed the low carrier densities, /cm3, and relatively high mobilities, typically cm2/V-s at helium temperature. Sharp bulk superconductivity was verified with both resistive transition and two-coils mutual inductance (MI) measurements. The scanning Tunneling Microscopy (STM) and spectroscopy (STS) measurements were performed after removing the top few surface layers with the STM tip itself. The details of this surface cleaning procedure are described in the Supplemental Material [33], and are based on a similar procedure previously used to expose superconducting regions in vacuum-cleaved YPtBi crystals [34].
@@
Fig. 1 shows representative base temperature (350 mK) spectra of pristine surface and tip-cleaned surface for YPtBi and LuPtBi samples. While the tip-cleaned surface improved the spectra, particularly in reducing the observed zero bias conductance to very low values, the pristine surface spectra pointed to a similar large gap. The observed spectra are rounded, even well below temperatures corresponding to the observed large gap. Such a behavior could be because of a nodal gap structure or high effective temperature due to RF heating. The observation of a flat gap (Fig. 1b) rules these options out. Another possibility is that a thin metallic layer between the superconducting state and the tip depresses the gap by proximity effect (Fig. 1c) [37]. In this scenario, the variability in the flatness of the gap and the density of states at zero bias [33] is explained by the inhomogeneous metallic film that is created by the tip cleaning. Figure 1c illustrates the proposed effect of the current pulse. Before the pulse, a thin disordered metallic layer covers the surface, and the pulse removes some, but not all, of this layer, resulting in a variable metallic barrier, which rounds or supresses the gap at different spots.
The large gaps observed at base temperature indicate that there is a superconducting system in both LuPtBi and YPtBi which involves a transition temperature higher than the bulk. Indeed, temperature dependence of tunneling spectra of both materials indicates gaps that persist much above the bulk . Fig. 2a shows a typical set of spectra taken on YPtBi at different temperatures. To estimate the gap, we fitted the spectra to the Dynes formula [38, 33]. The temperature was left as a fitting parameter, with a higher effective temperature required to account for the broadening of the spectra. Figure 2b shows the extracted gaps plotted against the measured temperature. The gap at our base temperature was 0.4 meV. Assuming yields . We note that while the spectra taken following the tip-cleaning procedure can be fitted to the Dynes formula, this procedure fails for the spectra taken on the pristine surface, due to the large residual density of states. This is consistent with the presence of a thin metallic barrier between the tip and the superconducting layer, as has been observed and modeled for thin metal-supercodnuctor interfaces [37]. This metallic barrier is then removed by the pulsing procedure.
The tip cleaning procedure could change the composition on the surface or the lattice structure. Indeed, the procedure used in Ref. [34] generated craters around the pulse location, in which a polycrystalline system was formed. The confinement of superconductivity to these craters could result in an increaed gap. However, the gaps we observed before and after the tip-cleaning procedure were of similar sizes, suggesting that the cleaning does not induce the larger gaps. Furthermore, the topography of our samples following the pulsing procedure did not show craters, and spectra did not systematically vary with the distance from the pulse [33].
@@
A tip-induced enhancement of superconductivity is further ruled out by measurements of the ac susceptibility of both materials, using a mutual inductance technique [40]. These crystals were not treated before the measurements. Both materials showed a large increase of the diamagnetic response at their respective bulk transition temperatures (Fig. 3), in agreement with previous reports [13, 14]. However, a careful measurement of YPtBi revealed a small increase, 0.5% of the change due to the bulk transition, which onsets at 3 K (Fig. 3a). LuPtBi showed a similar change, at around 5 K, over a background that is caused by thermal expansion (Fig. 3b and [33]). This may indicate that a small volume fraction of both samples turns superconducting at a higher temperature than the bulk.
@@
The observation of a similar transition temperature for the vanishing of the superconducting gap in tunneling, and for the onset of an inductive signal in mutual inductance points to an intrinsic origin of superconductivity, that onsets at a temperature much above the bulk . We envision two possible origins for the observations. First, given our inability to make clean, atomically flat cleaves that are known to expose the (111) plane, the signals may be attributed to another compound, created during the surface cleaning, a combination of the HH elements, possibly with contaminants, particularly oxygen. Superconducting compounds that may be formed this way are listed in Table 1. While the of both amorphous and polycrystalline Bi are suitable, such compounds on the surface would generate similar transitions for both LuPtBi and YPtBi, inconsistent with the difference in gap sizes and onset temperatures. We further note that LuPtBi and YPtBi are very similar chemically and structurally (lattice constants Å and Å, respectively [11]), suggesting that a strained inclusion of a different compound on the surfaces would yield a similar gap. Moreover, the similarity in the behavior of LuBi and YBi may rule this possibility out as well, particularly since they require pressure to exhibit superconductivity.
Compound | Bi | a-Bi | p-Bi | BiPt | Bi2Pt | YBi | LuBi |
@@
The alternative explanation for the observed higher in the two compounds is the creation of a chiral surface state. Particularly, since both LuPtBi and YPtBi are known to be topological with a large band inversion [11], it was suggested that superconductivity associated with this material is topological and may break TRS [22, 23]. To test this possibility, we performed high resolution polar Kerr effect measurements on YPtBi using a zero-area-loop Sagnac interferometer (ZALSI) [35, 36] (see supplementary material [33]).
@@
We cooled down the sample to 300 mK from a high temperature in the presence of a small external magnetic field, and then removed the field and measured the Kerr signal while warming up the sample (Fig. 4a). For small training fields ( Oe) the initial signal (at 300 mK) following the removal of the magnetic field was very small, consistent with zero (within our resolution of rad). However, increasing the temperature from base temperature, a finite Kerr signal of order rad appeared above the bulk . The signal was independent of the field strength and changed sign upon changing the direction of the training field (Fig. 4b and [33]). These features indicate a TRS breaking state, and are generally similar to those observed in chiral superconductors [49, 50]. Note that a TRSB state further restricts the association of the signals observed using mutual inductance and STM with a different chemical structure of the surface. In particular, amorphous and polycrystalline Bi are not chiral superconductors, and therefore will not give rise to a finite Kerr effect. The contribution from vortices [33] trapped by the field-cooling procedure is too small to account for the observed signals.
@@
With strong structural and electronic structure similarities to GdPtBi, we estimate the optical penetration depth of YPtBi at nm [51]. Thus, the Kerr signal can have both bulk and surface contributions. In particular above the bulk the observed Kerr signal may be a result of TRS breaking associated with the surface superconductivity - a favorable possibility since Kerr effect persists very close to the surface , or it may be associated with vestigial order [52, 53, 54, 55] originating from the bulk superconducting state. However, the suppression of the Kerr rotation below , as well as the nodes in the gap and analysis of the possible spin and orbital pairing states [23] suggest that the bulk does not break TRS.
@@
While low field training (i.e. training in Oe) yields similar onset and size of Kerr signal, which follows the direction of the training field, cooling the sample at higher magnetic field to base temperature (0.3 K), turning the field to zero and measuring while warming up the sample in zero magnetic field yields a signal that is zero within our resolution of rad (Fig. 5a). However, cooling the sample at higher magnetic field to base temperature (0.3 K), but on the way turning the magnetic field to zero at K, then measuring the Kerr effect at zero field from base temperature up yields results similar to the low field training experiments (Fig. 5b,c).
@@
Such a peculiar effect may point to a delicate interplay between the vortex state that appears in the bulk material upon cooling in a field much below the bulk , and the TRSB surface state. For example, cooling at lower field and removing the field at low temperatures may result in low vortex occupancy due to weak surface and geometrical barriers (note that Oe, while the thermodynamic critical field is Oe for the bulk superconducting state of these materials). Above a field of order Oe, enough vortices may remain in the bulk, which upon crossing the bulk Tc will move out of the sample, thus possibly “scrambling” the single domain nature of the TRS-breaking state of the surface superconductivity. Thus, the magnetization of the surface state interacts with a conventional bulk superconducting order parameter. This is similar to magnetic superconductors, where the formation and motion of vortices can change the ferromagnetic order parameter [56].
@@
Taken together, the STM, mutual inductance and PKE results indicate that the (111) surfaces of YPtBi and LuPtBi host a superconudcting state which breaks TRS. To understand the origin of this state, we studied the surface band structure of LuPtBi (see Supplementary Information [33]). In Bi-terminated (111) surfaces, we found a van-Hove signularity close to the -point, resulting in an enhanced density of states near the Fermi level, which could give rise to the enhanced surface . The calculated band structure of the Lu terminated surface did not show van-Hove singularities. Schwemmer et al. [7] have recently reported a detailed study of the surface band structure and possible superconductivity of the Bi-terminated (111) surface of LuPtBi. While the gross features of their surface band structure are similar to our calculation, they also revealed flat-band regions near the Fermi energy, and calculated the propensity to a superconducting instability. In particular, they argue that the transition temperature of the surface state is independent of the bulk transition, and exhibit vestigial time-reversal symmetry breaking in it fluctuation regime. Since inversion symmetry is broken, simple singlet and/or triplet states were dismissed and pairing was tested for pairs of electrons that are connected by the TR operator. The result has been a pair-amplitude that transforms under the E irreducible representation of the surface point group (relevant to the (111) crystal orientation) in the presence of strong Kane-Mele type and Rashba type spin-orbit coupling [57]. Furthermore, they found that a chiral superconducting state, which spontaneously breaks TRS, is energentically favorable. While we focused on the (111) oriented surface, surface bands close to the Fermi energy that originate from dangling bonds are generic in the half-Heuslers along other surfaces, including the (001) [12, 58] and (120) [30]. It is therefore possible that the surface superconductivity we report here could also exist on other surfaces, depending on the filling and flatness of the surface bands.
@@
In conclusion, we presented a comprehensive study of the (111) surface of the half-Heusler alloys YPtBi and LuPtBi. While a bulk of 0.8 K is observed for both systems, we find a much larger corresponding gap in our tunneling studies with ’s of 2.3 K and 6.5 K for YPtBi and LuPtBi, respectively. Mutual inductance measurements of YPtBi further revealed a small diamagnetic response that onsets around . Polar Kerr effect on YPtBi indicates that TRS is broken near or at . Together, our data suggest that chiral surface superconductivity precedes bulk condensation in these materials, a hypothesis which is strongly supported by recent calculations [7]. While additional work is required to determine the order parameter of these states, our results open a new direction to engineering topological superconducting states in half Heusler compounds. Although our study was conducted on single crystals, Kim et al. [59] recently demonstrated epitaxially growth of YPtBi films, which can also be encapsulated by an alumina film, preventing degradation of the surface. This system is therefore highly promising for quantum device applications.